OpenStudy (anonymous):

Find the inverse of the one-to-one function. f(x) = 2x + 5/7 Is this one y=7y+5/2

OpenStudy (anonymous):

oops that is suppose to be y=7x+5/2

OpenStudy (anonymous):

solve implicite for x and then replace x by y, this gives you the inverse function

OpenStudy (anonymous):

So is that what I did my getting y=7x + 5 /2

OpenStudy (anonymous):

\[ y=2x+\frac{5}{7} \\ \therefore x = \frac{1}{2}\left(y-\frac{5}{7}\right)\]

OpenStudy (anonymous):

\[ f(\bar{x})=\frac{1}{2}\left(x-\frac{5}{7}\right)\]

OpenStudy (anonymous):

but my choice look like this a. f^-1(x) = 7/2x+5 b. f^-1(x) = 7 / 2x -5 c. f^-1(x) = 7x + 5 /2 d. f^-1(x) = 7x - 5 / 2

OpenStudy (anonymous):

Sorry, but in this case I can't help you, since I don't find any of the solutions above representing a inverse function of the one you have posted

OpenStudy (anonymous):

See that is what is confusing me because the answer I got to didn't match my choices

OpenStudy (anonymous):

thanks any ways I will call the school tomorrow about this one

OpenStudy (anonymous):

alright, you're welcome

OpenStudy (anonymous):

Keep in mind that you can verify your answer always for yourself, by drawing it in a xy-Coordinate system, it should be mirrored at y=x, which is the case for the function we have.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

and also f(x) when you replace it by the inverse function for the original equation should be x, which is also the case for the function we have.